cos 35 degree/sin 35 degree + sin 11 degree/cos 79 degree - cos 28 degree cosec 62 degree
Answer by Alan3354(69443)

cos 35 degree/sin 35 degree + sin 11 degree/cos 79 degree - cos 28 degree cosec 62 degree
------------
If you mean:
cos(35)/sin(35) + sin(11)/cos(79) - cos(28)*csc(62)
Don't clutter it up with gibberish.
sin(x) = cos(90-x)
-------------------
cot(35) + sin(11)/sin(11) - sin(62)*(csc(62)
= cot(35) + 1 - 1
= cot(35)